# The Bayesian Origin of the Probability Weighting Function in Human Representation of Probabilities

### v1

@@ B.1.2 Properties of BLO Model
B.2 FI for Efficient Code in Frydman & Jin (2023)
Frydman & Jin (2023) take the encoding to be given by the code derived in Heng et al. (2020), which is defined via:
@@ Appendix C Experimental Methods and Supplementary Results
C.1 Details on Task and Datasets
C.1.1 Judgment of Relative Frequency (JRF) Task in Section 4.1
For this task, we analyzed the datasets used by Zhang et al. (2020), both publicly available. On each trial, subjects were briefly shown an array of black and white dots and reported their estimate of the relative frequency of one color by clicking on a horizontal scale.
@@ C.1 Details on Task and Datasets
  • Zhang et al. (2020) Dataset (JD, including JDA and JDB):
@@
In both datasets, the total number of dots on any given trial was one of five values: 200, 300, 400, 500, or 600. For all datasets, each trial included the true stimulus proportion, the subject’s estimate, total number of dots and other information, such as reaction time, that we didn’t use for model fitting.We did not apply any data preprocessing.
C.1.2 Pricing Task in Decision-making Under Risk (DMR) in Section 4.2
For this task, we analyzed the publicly available dataset from Zhang et al. (2020). This task used the same procedure and design as Gonzalez & Wu (1999). On each trial, subjects were presented with a two-outcome monetary gamble (e.g., a 50% chance to win \$100 or \$0 otherwise) and a table of sure amounts. subjects made a series of choices between the gamble and the sure amounts, a process that used a sequential bisection method to narrow the range and determine their Certainty Equivalent (CE) for the gamble.
@@
C.1.3 Choice Task in Decision-making Under Risk (DMR) in Section 4.2
@@
C.2 Details for Adaptation Experiment in Section 4.3
This experiment used the JRF dot-counting paradigm, but critically, the distribution of stimulus proportions was manipulated to be bimodal to test the model’s ability to adapt its prior. Data was collected on an online platform following the procedure in Zhang et al. (2020).
@@
For all subjects in our collected datasets, in addition to the true number of black and white dots, the color designated for estimation, the subject’s estimated proportion, and the reaction time, we also varied and logged the display time for each trial.
We removed estimated proportions of 0 and 100, as these would have zero likelihood under the BLO model.
C.3 Fitting models to Datasets
Across all datasets, we ran all Bayesian models using the implementation from Hahn & Wei (2024); Hahn et al. (2025), using a grid of 200 points and regularization strength 1.0.
For all the Bayesian models, as well as the reimplemented Zhang et al. (2020)’s models, the optimization was performed using a gradient-based approach(Adam or SignSGD).
C.4 Details on Model Fit Metrics
@@
This is the metric used in Zhang et al. (2020). This metric assesses the overall quality of a model by balancing its performance with its simplicity. The procedure of using this metric is as follows:For each subject, a model is trained on their entire dataset. This follows the procedure used in the work you are comparing against.
@@
    For each subject, a model is trained on all the trials. This follows the procedure used Zhang et al. (2020). We get the model’s final NLL and count its number of free parameters ($k$).
@@
    We then calculate the Corrected Akaike Information Criterion (AICc) using the final NLL and $k$. The AICc formula essentially adds a penalty to the NLL for each extra parameter: $\text{AICc}\approx 2\cdot\text{NLL}+2k$.
@@
C.5 Details on Categorical and Analytical Fitting
@@ C.5.2 DMR Pricing Task
In Zhang et al. (2020)’s dataset, $\text{CE}_{\max}=800$ and $\text{CE}_{\min}=0$. We use 1000 grid size for the categorical version.
C.6 Details on Bayesian Model for JRF Task
C.6.1 Bayesian Model Variants
@@
C.6.2 Performance Comparison of All Model variants
@@
Overall, our model performance better than Zhang et al. (2020)’s model variants across two metrics and two fitting approaches. Within Bayesian models, models using a bounded log-odds encoding outperform those with an unbounded encoding, and a Gaussian prior is superior to a uniform prior. Within Zhang et al. (2020)’s model variants, bounded model performs better than unbounded model, and using V(p)V(p)(explained in Eq 39) is better than assuming a constant value $V$.
@@ B Categorical
C.6.3 Analysis of Fitted Prior and Resources
@@
Figure 10: JRF Task: Per-subject resources ($\sqrt{\mathcal{J}(p)}$) for the Bayesian model with freely fitted prior(in red) and encoding and the BLO model(in blue). Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from Zhang et al. (2020)), and S76 to S86 from dataset ZM12 (Zhang & Maloney (2012)).
Figure 11: JRF Task: Discontinuity observed in the resources ($\sqrt{\mathcal{J}(p)}$) of the BLO model. The proof of this property is in Theorem 8. Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from Zhang et al. (2020)), and S76 to S86 from dataset ZM12 (Zhang & Maloney (2012)).
@@
Figure 13: JRF Task: Group-level resources ($\sqrt{\mathcal{J}(p)}$) for the four encoding components evaluated in this task, along with the resources of the BLO model. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.13. This result is mentioned in main text section 4.1.
C.6.4 Analysis of Bias and Variance
@@
Figures 15 and 16 show the per-subject bias and variance, respectively, providing a more detailed view of the group-level results presented in the main text (Figure 3). The methods used to calculate these quantities are detailed in Appendix C.14.
@@
Figure 14: JRF Task: The decomposition of the bias into attraction and repulsion. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.13.
Figure 15: JRF Task: Per-subject bias of the non-parametric estimates from the data, the BLO model and the Bayesian model with free prior and free encoding. Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from Zhang et al. (2020)), and S76 to S86 from dataset ZM12 (Zhang & Maloney (2012)).
Figure 16: JRF Task: Per-subject variability of the non-parametric estimates from the data, the BLO model and the Bayesian model with free prior and free encoding. Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from Zhang et al. (2020)), and S76 to S86 from dataset ZM12 (Zhang & Maloney (2012)).
C.7 Details on Bayesian Model for DMR Pricing Task
C.7.1 Bayesian Model Variants and Fitting Procedure
@@
The goal of this initial stage was to convert each raw CE response into a non-parametric, trial-level estimate of the subjective probability weight, $\hat{\pi}_{\text{implied},t}(p)$. For each trial $t$, we fits a free variable. We also fit $\alpha$ for applying the CPT utility function. To account for additional variability in CE, we included an extra noise term $\epsilon_{\text{CE}}$ and optimized parameters by minimizing the loss between predicted and observed CE. This method is similar to that of Zhang et al. (2020), but our implementation works on a trial-by-trial basis rather than on 11 discrete probability levels.
@@
C.7.2 Performance Comparison of All Model Variants
@@ B Categorical
C.7.3 Analysis of Fitted Prior and Resources
@@
Figure 19: DMR Pricing Task: Group-level resources ($\sqrt{\mathcal{J}(p)}$) for the three encoding components evaluated in this task. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.13.
@@
C.7.4 Analysis of Bias and Variance
Figure 21: DMR Pricing Task: Per-subject bias of the non-parametric estimates from the data, the BLO model and the Bayesian model with free prior and free encoding on the DMR pricing task. Subjects S1 to S51 are from dataset JDA, and S52 to S75 are from dataset JDB (both from Zhang et al. (2020)).
@@ C.7.4 Analysis of Bias and Variance
C.8 Details on Bayesian Model for DMR Choice Task
C.8.1 Full Utility Function and Logit Choice Rule
@@
The parameter $\tau$ ($\tau>0$) controls the choice sensitivity and is a free parameter to fit.
C.8.2 Bayesian Model Variants and Fitting Procedure
@@
C.8.3 Performance Comparison of All Model Variants
@@
C.8.4 Analysis of Fitted Prior and Resources
@@
Figure 23: DMR Choice Task: Group-level resources ($\sqrt{\mathcal{J}(p)}$) for the four encoding components evaluated in this task. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.13.
@@ C.8.5 Analysis of Bias
Figure 25: DMR Choice Task: Per-subject bias of the LILO model, the Prelec model and the Bayesian model with free prior and free encoding on the DMR pricing task. Subjects S1 to 153 are from dataset CPC15 (Erev et al., 2017).
C.9 Details on Bayesian Model for Adaptation Bimodal Data
C.9.1 Bayesian Model Variants
@@
C.9.2 Performance Comparison of All Model variants
@@ B Categorical
C.9.3 Analysis of Fitted Prior and Resources
@@
Figure 28: Adaptation Task: Group-level resources ($\sqrt{\mathcal{J}(p)}$) for the three encoding components evaluated in this task. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.13.
@@
C.9.4 Analysis of Bias and Variance
@@
Figure 32: Adaptation Task: Bias under bimodal stimulus statistics. Each panel shows estimation bias as a function of $\Delta p$ (stimulus proportion relative to the trough between the two stimulus peaks). Thin curves show individual subjects or model fits; thick curves show the across-subject or across-fit median. (Left) Human data: subjects exhibit multi-peaked biases, with attraction toward both stimulus modes. (Second) Bayesian model with freely estimated prior and encoding (FreeP, FreeE) successfully reproduces this multi-peaked structure. (Third) Bayesian model with FreeP but encoding constrained to match the prior (PriorMatchedE) predicts biases away from the prior modes, consistent with efficient coding but inconsistent with human data; it cannot capture the human pattern. (Right) BLO model fails to capture the human pattern. This figure is discussed in the main text section 4.3
C.10 Details on BLO Model for JRF Task and Adaptation Data
C.10.1 Application of BLO Model to JRF Task
@@
, where $N$ is the total number of dots and $n_{s}$ is a free parameter. We note that sampling of the items is not implemented in the published implementation from Zhang et al. (2020), and we correspondingly do not attempt to model it in our reimplementation of the BLO model.
C.10.2 Reimplementation of BLO and LLO Models Variants
We reimplemented the key model variants proposed by Zhang et al. (2020) for JRF. Specifically, we focused on four variants, Bounded Log-Odds(BLO) and Linear Log-Odds (LLO) models, each combined with either a constant perceptual variance (V=Const) or a proportion-dependent variance compensation (V=V(p)).
Following Zhang et al. (2020), we fit these models to the data per subject by minimizing the negative log-likelihood of the observed responses given the model’s predicitions. Our reimplementation achieved nearly identical estimated parameter values and model fits to those reported by Zhang et al. (2020) for JRF.
For all these reimplemented models, the optimization was performed using a gradient-based approach(Adam or SignSGD) rather than the original Nelder-Mead optimizer(fminsearchbnd) used in their implmentation. We set parameter bounds for our optimization based on Zhang et al. (2020)’s settings. The hyperbolic tangent (tanh) function was applied to parameterize model parameters such as $\Delta_{-}$ and $\Delta_{+}$, while the exponential function was used for strictly positive parameters like $\kappa$. Parameters not naturally constrained were optimized directly.
We successfully achieved nearly identical estimated parameter values and model fits to those reported by Zhang et al. (2020) for the dot counting task, thereby validating our reimplementation.
C.11 Details for BLO on DMR Pricing Task
C.11.1 Application of BLO Model to DMR Task

### v2

@@ B.1.2 Properties of BLO Model
B.2 FI for Efficient Code in frydman2023source
frydman2023source take the encoding to be given by the code derived in heng2020efficient, which is defined via:
@@ Appendix C Experimental Methods and Supplementary Results
C.1 Theoretical Mapping of Model Variants
Table 1: Correspondence between existing theoretical accounts and the specific model variants used in our empirical comparison.
Concept Corresponding Model in Our Comparison
Fixed probability distortion function LILO, Prelec
Log-odds based accounts BLO
Regression based accounts Bayesian model with uniform encoding
Efficient-coding Bayesian model with prior-matched encoding
General Bayesian account Bayesian model with freely fitted prior and encoding
C.2 Details on Task and Datasets
C.2.1 Judgment of Relative Frequency (JRF) Task in Section 4.1
For this task, we analyzed the publicly available dataset from zhang2020 and used ZM2012 . On each trial, subjects were briefly shown an array of black and white dots and reported their estimate of the relative frequency of one color by clicking on a horizontal scale.
@@ C.2 Details on Task and Datasets
  • zhang2020 Dataset (JD, including JDA and JDB):
@@
In both dataset, the total number of dots on any given trial was one of five values: 200, 300, 400, 500, or 600. For all datasets, each trial included the true stimulus proportion, the subject’s estimate, total number of dots and other information, such as reaction time, that we didn’t use for model fitting.We did not apply any data preprocessing.
C.2.2 Pricing Task in Decision-making Under Risk (DMR) in Section 4.2
For this task, we analyzed the publicly available dataset from zhang2020. This task used the same procedure and design as gonzalez1999shape. On each trial, subjects were presented with a two-outcome monetary gamble (e.g., a 50% chance to win \$100 or \$0 otherwise) and a table of sure amounts. subjects made a series of choices between the gamble and the sure amounts, a process that used a sequential bisection method to narrow the range and determine their Certainty Equivalent (CE) for the gamble.
@@
C.2.3 Choice Task in Decision-making Under Risk (DMR) in Section 4.2
@@
C.3 Details for Adaptation Experiment in Section 4.3
This experiment used the JRF dot-counting paradigm, but critically, the distribution of stimulus proportions was manipulated to be bimodal to test the model’s ability to adapt its prior. Data was collected on an online platform following the procedure in zhang2020.
@@
For all subjects in our collected datasets, in addition to the true number of black and white dots, the color designated for estimation, the subject’s estimated proportion, and the reaction time, we also varies and logged the display time for each trial.
We removed estimated proportions of 0 and 100.
C.4 Fitting models to Datasets
Across all datasets, we ran all Bayesian models using the implementation from hahn2024unifying; Hahn2025Identifiability, using a grid of 200 points and regularization strength 1.0.
For all the Bayesian models, as well as the reimplemented zhang2020’s models, the optimization was performed using a gradient-based approach(Adam or SignSGD).
C.5 Details on Model Fit Metrics
@@
This is the metric used in zhang2020. This metric assesses the overall quality of a model by balancing its performance with its simplicity. The procedure of using this metric is as follows:For each subject, a model is trained on their entire dataset. This follows the procedure used in the work you are comparing against.
@@
    For each subject, a model is trained on all the trials. This follows the procedure used zhang2020. We get the model’s final NLL and count its number of free parameters ($k$).
@@
    We then calculate the Corrected Akaike Information Criterion (AICc) using the final NLL and $k$. The AICc formula essentially adds a penalty to the NLL for each extra parameter: $\text{AICc}\approx 2*\text{NLL}+2k$.
@@
For the nonparametric Bayesian models (e.g., FreeE, FreeP), defining the effective number of parameters is non-trivial due to regularization, which reduces the effective degrees of freedom below the raw grid size. To provide a lower bound on the model performance, we computed a maximally conservative AICc using the upper bound of model parameters. For example, we use 403 as the number of parameters of FreeE, FreeP model in JRF task(200 for encoding, 200 for prior, and 3 noise or mixing parameters). Because this assumption over-penalizes the nonparametric models, we do not include these values in the comparison bar plots. Instead, we report them in the corresponding footnotes.
C.6 Details on Categorical and Analytical Fitting
@@ C.6.2 DMR Pricing Task
In zhang2020’s dataset, $\text{CE}_{\max}=800$ and $\text{CE}_{\min}=0$. We use 1000 grid size for the categorical version.
C.7 Details on Bayesian Model for JRF Task
C.7.1 Bayesian Model Variants
@@
C.7.2 Performance Comparison of All Model variants
@@
Overall, our model performance better than zhang2020’s model variants across two metrics and two fitting approaches. Within Bayesian models, models using a bounded log-odds encoding outperform those with an unbounded encoding, and a Gaussian prior is superior to a uniform prior. Within zhang2020’s model variants, bounded model performs better than unbounded model, and using V(p)V(p)(explained in Eq 41) is better than assuming a constant value $V$.
@@ B Categorical
C.7.3 Analysis of Fitted Prior and Resources
@@
Figure 10: JRF Task: Per-subject resources ($\sqrt{J(p)}$) for the Bayesian model with freely fitted prior(in red) and encoding and the BLO model(in blue). Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from zhang2020), and S76 to S86 from dataset ZM12 (ZM2012).
Figure 11: JRF Task: Discontinuity observed in the resources ($\sqrt{J(p)}$) of the BLO model. The proof of this property is in Theorem 8. Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from zhang2020), and S76 to S86 from dataset ZM12 (ZM2012).
@@
Figure 13: JRF Task: Group-level resources ($\sqrt{J(p)}$) for the four encoding components evaluated in this task, along with the resources of the BLO model. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14. This result is mentioned in main text section 4.1.
C.7.4 Analysis of Bias and Variance
@@
Figures 14 and 15 show the per-subject bias and variance, respectively, providing a more detailed view of the group-level results presented in the main text (Figure 3). The methods used to calculate these quantities are detailed in Appendix C.15.
@@
Figure 14: JRF Task: Per-subject bias of the non-parametric estimates from the data, the BLO model and the Bayesian model with free prior and free encoding. Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from zhang2020), and S76 to S86 from dataset ZM12 (ZM2012).
Figure 15: JRF Task: Per-subject variability of the non-parametric estimates from the data, the BLO model and the Bayesian model with free prior and free encoding. Subjects S1 to S51 are from dataset JDA, S52 to S75 from dataset JDB (both from zhang2020), and S76 to S86 from dataset ZM12 (ZM2012).
C.8 Details on Bayesian Model for DMR Pricing Task
C.8.1 Bayesian Model Variants and Fitting Procdure
@@
The goal of this initial stage was to convert each raw CE response into a non-parametric, trial-level estimate of the subjective probability weight, $\hat{\pi}_{\text{implied},t}(p)$. For each trial $t$, we fits a free variable. We also fit $\alpha$ for applying the CPT utility function. To account for additional variability in CE, we included an extra noise term $\epsilon_{\text{CE}}$ and optimized parameters by minimizing the loss between predicted and observed CE. This method is similar to that of zhang2020, but our implementation works on a trial-by-trial basis rather than on 11 discrete probability levels.
@@
C.8.2 Performance Comparison of All Model Variants
@@ B Categorical
C.8.3 Analysis of Fitted Prior and Resources
@@
Figure 18: DMR Pricing Task: Group-level resources ($\sqrt{J(p)}$) for the three encoding components evaluated in this task. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14.
@@
C.8.4 Analysis of Bias and Variance
Figure 20: DMR Pricing Task: Per-subject bias of the non-parametric estimates from the data, the BLO model and the Bayesian model with free prior and free encoding on the DMR pricing task. Subjects S1 to S51 are from dataset JDA, and S52 to S75 are from dataset JDB (both from zhang2020).
@@ C.8.4 Analysis of Bias and Variance
C.9 Details on Bayesian Model for DMR Choice Task
C.9.1 Full Utility Function and Logit Choice Rule
@@
The parameter $\tau$ ($\tau>0$,) controls the choice sensitivity and is a free parameter to fit.
C.9.2 Bayesian Model Variants and Fitting Procedure
@@
C.9.3 Performance Comparison of All Model Variants
@@
[…]
C.9.4 Analysis of Fitted Prior and Resources
@@
Figure 22: DMR Choice Task: Group-level resources ($\sqrt{J(p)}$) for the four encoding components evaluated in this task. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14.
@@ C.9.5 Analysis of Bias
Figure 24: DMR Choice Task: Per-subject bias of the LILO model, the Prelec model and the Bayesian model with free prior and free encoding on the DMR pricing task. Subjects S1 to 153 are from dataset CPC15 (erev2017anomalies).
C.10 Details on Bayesian Model for Adaptation Bimodal Data
C.10.1 Bayesian Model Variants
@@
C.10.2 Performance Comparison of All Model variants
@@ B Categorical
C.10.3 Robustness of Model Comparison via Bootstrapping
[…]
Table LABEL:tab:bootstrap summarizes the results. We report the Bootstrap Support, the percentage of resampled datasets in which the FreeP, FreeE model achieved a lower Summed NLL than the competitor. We also report the 95% Confidence Interval (CI) of the difference in NLL (competing model minus FreeP, FreeE).
The freely fitted Bayesian model significantly outperformed all competing accounts, including the Efficient Coding model (PriorMatchedE) and the BLO model, achieving a bootstrap support greater than $98\%$. The 95% confidence intervals for the NLL differences exclude zero for all comparisons. This analysis confirms the superiority of our Bayesian account in this task.
[…]
FreeP, PriorMatchedE (Efficient Coding) 98.1% $[233.43,7550.74]$
BimodalP, BoundedLOE 100.0% $[7177.88,13032.97]$
BimodalP, UnboundedE 100.0% $[7216.18,13054.51]$
GaussianP, BoundedLOE 100.0% $[7380.49,13224.09]$
(BLO) Bounded, $V=V(p)$ 100.0% $[7518.55,13399.81]$
GaussianP, UnboundedE 100.0% $[7466.34,13271.41]$
Bounded, $V=\text{Const}$ 100.0% $[7532.97,13426.86]$
Unbounded, $V=V(p)$ 100.0% $[7518.60,13437.80]$
Unbounded, $V=\text{Const}$ 100.0% $[7510.38,13429.20]$
C.10.4 Analysis of Fitted Prior and Resources
@@
Figure 27: Adaptation Task: Group-level resources ($\sqrt{J(p)}$) for the three encoding components evaluated in this task. The solid line shows the group median, and the shaded area indicates the interquartile range (IQR). Plotting details are provided in Appendix C.14.
@@
C.10.5 Analysis of Bias and Variance
@@
Figure 31: Adaptation Task: Bias under bimodal stimulus statistics. Each panel shows estimation bias as a function of $\Delta p$ (stimulus proportion relative to the trough between the two stimulus peaks). Thin curves show individual subjects or model fits; thick curves show the across-subject or across-fit median. (Left) Human data: subjects exhibit multi-peaked biases, with attraction toward both stimulus modes. (Second) Bayesian model with freely estimated prior and encoding (FreeP, FreeE) successfully reproduces this multi-peaked structure. (Third) Bayesian model with FreeP but encoding constrained to match the prior (PriorMatchedE) predicts biases away from the prior modes, consistent with efficient coding but inconsistent with human data. (Right) BLO model fails to capture the human pattern. This figure is dicussed in the main text section 4.3
C.11 Details on BLO Model for JRF Task and Adapatation Data
C.11.1 Application of BLO Model to JRF Task
@@
, where $N$ is the total number of dots and $n_{s}$ is a free parameter. We note that sampling of the items is not implemented in the published implementation from zhang2020, and we correspondingly do not attempt to model it in our reimplementation of the BLO model.
C.11.2 Reimplementation of BLO and LLO Models Variants
We reimplemented the key model variants proposed by zhang2020 for JRF. Specifically, we focused on four variants, Bounded Log-Odds(BLO) and Linear Log-Odds (LLO) models, each combined with either a constant perceptual variance (V=Const) or a proportion-dependent variance compensation (V=V(p)).
Following zhang2020, we fit these models to the data per subject by minimizing the negative log-likelihood of the observed responses given the model’s predicitions. Our reimplementation achieved nearly identical estimated parameter values and model fits to those reported by zhang2020 for JRF.
For all these reimplemented models, the optimization was performed using a gradient-based approach(Adam or SignSGD) rather than the original Nelder-Mead optimizer(fminsearchbnd) used in their implmentation. We set parameter bounds for our optimization based on zhang2020’s settings. the hyperbolic tangent (tanh) function was applied to parameters such as $\Delta^{-}$ and $\Delta^{+}$ , while the exponential function was used for strictly positive parameters like $\kappa$. Parameters not naturally constrained were optimized directly.
We successfully achieved nearly identical estimated parameter values and model fits to those reported by zhang2020 for the dot counting task, thereby validating our reimplementation.
C.12 Details for BLO on DMR Pricing Task
C.12.1 Application of BLO Model to DMR Task